Denote by 𝑄𝑅 and 𝑆 the projections of the point P=(−2,3,−4) onto the 𝑥𝑦 plane, the 𝑦𝑧 plane, and the 𝑥𝑧 plane, respectively. Which of the following line segments has the greatest length?


1
Expert's answer
2021-02-25T05:08:42-0500

Solution:

We know that "xy" -plane is described by equation "z=0" .

So, projection of "P(-2,3,-4)" on "xy" -plane is "Q(-2,3,0)" .

We know that "yz" -plane is described by equation "x=0" .

So, projection of "P(-2,3,-4)" on "yz" -plane is "R(0,3,-4)" .

We know that "xz" -plane is described by equation "y=0" .

So, projection of "P(-2,3,-4)" on "xz" -plane is "S(-2,0,-4)" .

Now, we find their distances or length with "O(0,0,0)" using distance formula.

"QO=\sqrt{(-2-0)^2+(3-0)^2+(0-0)^2}=\sqrt{4+9}=\sqrt{13}"

"RO=\sqrt{(0-0)^2+(3-0)^2+(-4-0)^2}=\sqrt{9+16}=\sqrt{25}"

"SO=\sqrt{(-2-0)^2+(0-0)^2+(-4-0)^2}=\sqrt{4+16}=\sqrt{20}"

Clearly, "\sqrt{25}" is greater, thus RO is the longest length.